Print

Print


Ticket Number: HEA-NOC/20080205-8      Ticket Status: CLOSE
Ticket Type:   scheduled               Resolver:      HEA-NOC
Ticket Opened: 20080205 12:05 UTC      Problem Start: 20080205 09:00 UTC
Ticket Closed: 20080219 13:19 UTC      Problem End:   20080219 13:18 UTC

Site/line:     DNS, Listserv, Multicast, NDLR, Netflow, Web Hosting,
               eduroam

Problem Description:
Package updates are required on the following machines:


carnegie
howth
jesse
lapsang
mellon
misha
porthos
samhain

Problem Effects:
As these updates install a new kernel
(http://secunia.com/advisories/28706/), a reboot of each machine and
downtime of associated services will be required.


Problem History:
20080205 12:03 RG Open ticket to track work.

20080206 10:06 LT lapsang (aka ietlr2.heanet.ie, one of the resilient
national eduroam gateways) upgraded and rebooted. No impact on eduroam
service as traffic continued to be handled by the other resilient gateway.

20080211 11:19 RG howth, carnegie and misha rebooted.

20080213 09:47 RG jesse upgraded.

20080215 14:46 RG Mellon rebooted and upgraded.

20080218 18:10 GAR samhain failed to come back correctly after reboot.
halt and shutdown commands failed to restart system. Power cycled using
PDU. Currently working on restoring service to samhain. Until machine is
recovered, some web services will be unavailable.

20080219 13:18 GAR Closing ticket as further Debian updates will be
applied in future tickets.

samhain restored to service and under monitoring.

Time to Fix:
1 hour

-- 
This ticket can be monitored at http://www.hea.net/tickets/20080205-8

HEAnet Limited. Registered in Ireland, No. 275301.